marcy1030:

Meg skateboards mc003-1.jpg of a mile in mc003-2.jpg of an hour. At this rate, how far will she skateboard in an hour?

Vocaloid:

can you post a screenshot? I cannot access these file names

Vocaloid:

or type what numbers are?

marcy1030:

A motorist is pumping gas into his car at a rate of mc013-1.jpg of a gallon every mc013-2.jpg of a minute. At this rate, how many gallons of gas will he have pumped into his car in mc013-3.jpg of a minute?

Vocaloid:

I still cannot see what the numbers are (mc013-1.jpg) and (mc013-2.jpg) please type out what you see on the screen

marcy1030:

ok

marcy1030:

5/12

marcy1030:

1/24

marcy1030:

1/2

Vocaloid:

first you want to divide gallons/minutes to get a gallon per minute rate

Vocaloid:

(5/12)/(1/24) = ?

marcy1030:

i dont know it do u?

Vocaloid:

to divide by a fraction, take the reciprocal of the second fraction and multiply

Vocaloid:

(5/12)/(1/24) = (5/12) * (24) = ?

marcy1030:

10

Vocaloid:

awesome now, to calculate the number of gallons pumped in 1/2 a minute we would simply multiply 10 * (1/2) = ?

marcy1030:

5

Vocaloid:

good so the answer is 5 gallons

Vocaloid:

would you like to go over the skateboard problem?

Vocaloid:

I can't see what the numbers are but you would just divide the first number (distance) by the second number (time)

marcy1030:

Meg skateboards mc003-1.jpg of a mile in mc003-2.jpg of an hour. At this rate, how far will she skateboard in an hour?

marcy1030:

1/4

marcy1030:

1/2

marcy1030:

hello

Vocaloid:

sorry

Vocaloid:

just divide (1/4)/(1/2)

Ultrilliam:

Welcome to QC!